# Transfer matrix approach to quantum systems subject to certain Lindblad evolution

### v1

@@ III Transfer Matrix Approach
$\begin{pmatrix}a_{1}(q)&&a_{2}(q)&&a_{3}(q)&&a_{4}(q)\\ 1&&0&&0&&0\\ 0&&1&&0&&0\\ 0&&0&&1&&0\\ \end{pmatrix},$
@@
Once the transfer matrix is known, behavior of the product can be studied numerically, in case simple analytical solutions aren’t possible, for large and small $s$ limit for thermodynamically large system sizes, which as we will see in the next section determines the asymptotic behavior of the correlators at short and long times, respectively.
We conclude this section by obtaining the solution for the off-diagonal elements of $C(t)$ in terms of $\mathcal{G}_{0,0}$. The $l^{th}$ off-diagonal elements can be written as,
@@
$\mathcal{G}_{l,0}=\frac{-1}{i\omega\sin z}\left[\sin(zl)+(i\omega\sin[z(l-1)]-s\sin zl)\mathcal{G}_{0,0}\right].$ (35)
@@ IV Thermodynamic limit and Asymptotic behavior
In the last section, we showed that the solution for all the correlators is basically determined by a single function $\mathcal{G}_{0,0}$ whose dependence on $L$ is explicitly given in Eq. (30). So, taking the thermodynamic limit is taking $L\rightarrow\infty$ for this function. In this limit, we have $\sin(zL)\approx-\frac{e^{-izL}}{2i}$ since $z=\arccos(-iu)$ has a positive imaginary part as $u>0$. Substituting $\sin(zL)\approx-\frac{e^{-izL}}{2i}$ in Eq. (30) $\mathcal{G}_{0,0}$ simplifies to,
@@
$=\frac{1}{\tilde{s}-4\gamma-i\omega(q)e^{iz}}=\frac{1}{\sqrt{\tilde{s}^{2}+\omega(q)^{2}}-4\gamma},$ (37)
@@
$z=\arccos(-iu)=\frac{\pi}{2}+i\log\big{[}u+\sqrt{u^{2}+1}\big{]}~{}.$ (38)
Note that expanding $\mathcal{F}(s,q)$ in the long wavelength limit i.e. around $q=0$, we get
@@
which is the Green’s function for the telegrapher’s equation, Eq. (3), with $v^{2}=8J^{2}$, as expected.
The solution for diagonal elements in the limit $L\rightarrow\infty$ reads,
where we have converted the sum over $q_{n}$ into an integral. Note that the thermodynamic limit naturally came out of the solution in Eq. (30), this is one advantage of using this method as opposed taking the thermodynamic limit via the spectral decomposition of $\mathcal{A}$, i.e. Eq. (14). Eq. (40) provides a simple and useful expression for the density for following reasons. The long time and short time behavior follow very elegantly by simply looking at this equation in small $s$ and large $s$ limit, respectively. Let us look at these two limits separately and for simplicity we consider an initial state such that, $c_{x}=\delta_{x,0}$ which gives $c(q)=1$.
Short time limit: In this limit, we consider $s\sim a(q)>4\gamma$, so dropping $4\gamma$ in the denominator of Eq. (40) we get,
@@
where the Bessel function, $J_{0}[\omega(q)t]$ arises from the Inverse Laplace of $1/\sqrt{s^{2}+\omega^{2}(q)}$, and the square of the Bessel $J_{x}(4t)$, known to be the Green’s function in absence of dephasing, comes by the subsequent integration over $q$. We see that the ballistic behavior starts decaying exponentially as soon as the dephasing is turned on.
@@
For the long time limit, we consider $4\gamma>s\sim\omega^{2}(q)$. Also, since the behavior is dominated by small $q$ i.e. long wavelengths, $\omega(q)\sim 4q$. Using this and extending the limits in the $q$ integral from $\infty$- ∞ to $\infty$ as the dominant contribution comes near $q=0$, we have
@@
where $D=2J^{2}/\gamma$ is the diffusion constant which is in accordance with the expectations from the telegrapher’s equation. Note that the small $s$ and $q$ behavior of $\mathcal{G}_{0,0}$ is given by $(s+(2J^{2}q^{2}/\gamma))^{-1}$, and its Laplace inverse gives the diffusion kernel i.e. $e^{-2J^{2}q^{2}t/\gamma}$, as expected. A deviation from this behavior in $\mathcal{F}(s,q)$ would give rise to anomalous behaviors.
@@
The integrand has branch points at $b_{\pm}=\pm i\omega(q)$, and we take the branch cut to run along the imaginary axis between the two branch points. There are also two poles located at $s_{\pm}=\pm\sqrt{16\gamma^{2}-\omega^{2}(q)}$. The poles lie on the real axis for $4\gamma>\omega(q)$ and on the branch cut along imaginary axis for $4\gamma<\omega(q)$. The branch cut joins the two Riemann sheets and the contours of integration, see Fig. 1, for the Laplace inverse are taken on the principal Riemann sheet.
The contribution due to the branch cut is ballistic, and is sensitive to the details of the spectrum of the Hamiltonian i.e. the dispersion given by $\omega(q)$. When the poles lie on the imaginary line, their contribution is ballistic akin to forward and backward propagating plane waves of frequency $\sqrt{\omega^{2}(q)-16\gamma^{2}}$. However, as $q$ changes such that $|\omega(q)|$ approaches $4\gamma$ the poles sweep along the imaginary axis towards the origin and meet at the origin at $4\gamma=|\omega(q)|$. For $4\gamma>|\omega(q)|$ the poles are off the branch cut and shift to the real axis but on different Riemann sheets joined by the branch cut. Therefore, one has to be careful about which pole contributes. $s_{+}$ shifts to the principal Riemann sheet, and $s_{-}$ shifts to the second Riemann sheet. Thus, only $s_{+}$ contributes and gives a diffusive contribution. This is also evident from Eq. (44) where the contribution from the the pole at $s=-2q^{2}/\gamma$, which is just $s_{+}$ at small $q$ shifted by $4\gamma$, gives rise to the diffusion kernel. Physically, the contribution of the pole at $s_{-}$ and $s_{+}$ correspond to evolution backward and forward in time, receptively. Because the diffusion equation (holding in small $q$ or long-wavelength limit) has a meaning only evolving forward in time (going backwards in time any inhomogeneity would get sharper, instead of smoother, violating relaxation and the 2nd law) only $s_{+}$ has to be kept. For the ballistic short-wavelength part at $4\gamma<\omega(q)$, where the dephasing scattering length $\sim 1/\gamma$ is larger than the wavelength $\sim 1/q$ and one deals with the wave equation. Therefore, both forward and backward in time evolution are physical and both contributions have to be included.
@@
Eq. (40) can be directly evaluated numerically by summing over the allowed values of $q$ in place of the integral over $q$ and taking the Laplace inverse numerically using standard approaches such as Talbot’s method Dingfelder and Weideman (2015); Cohen (2007) etc. A nice implementation of these methods can be found in mpmath mpmath development team (2023) python library. While for the XX model with dephasing we can plug in Eq. (40) into the library and get the density but, in general, one needs to multiply the transfer matrices and invert the product to get $\mathcal{G}_{0,0}$ which is then Laplace inverted for every allowed value of $q$. Therefore, assuming that the transfer matrices are of dimension $r$, the complexity of such a calculation scales as $L^{2}r^{6}$. Given that $r$ is of the order of range of couplings which is usually small, the complexity of this method is better than obtaining the density at any time $t$ via exact diagonalization or iterations of matrix Eq. (10) of size $\sim L^{2}$, which have a complexity of about $L^{6}$. Using the transfer matrix approach magnetization profiles for system sizes of the order of $10^{6}$ can be easily obtained which is much bigger than $L\approx 10^{4}$ feasible through diagonalization Varma et al. (2017).
Fig. 2 shows the evolution of the magnetization density, $\langle\sigma_{3}^{x}\rangle=-C_{x,x}(t)$ for $L=10^{5}$ starting from a domain wall initial state using Eq. (40). By domain wall initial state, we mean a product state of Eq. (16) where spins with $x<L/2$ are in up state and spins with $x>L/2$ are in the down state. The slow down of the propagation of the front is visibly apparent from Fig. 2 as $\gamma t$ crosses unit value.
@@
The plot for $\beta(t)$ is shown in panel (b) of Fig. 3. We find that $\beta(t)$ shows oscillations which decay with time, see the brown dots $(L=10^{5},\gamma=0.01)$ and red crosses ($L=1600$, $\gamma=0.01$) in Fig. 3b. The oscillations arise as at small times the Green’s function is given by Bessel functions, and their amplitude goes down as a power-law (nearly ${}1/t$). The bare time scale for the oscillations is given by $t\sim 1/J$, which in the scaled time, $\gamma t$, goes as $\gamma/J$. Therefore, as $\gamma$ is lowered the oscillations shift towards smaller and smaller values of $\gamma t$. We can see this in the blue dashed curve ($L=10^{5}$, $\gamma=0.001$) where the oscillations have already become too small to be captured by Talbot’s numerical Laplace inverse, and it gives a monotonically decreasing curve from ballistic behavior, $\beta=1$, to diffusive behavior, $\beta=0.5$. Talbot’s method fails to capture the oscillations as the oscillation strength becomes small because of its limitations with oscillatory time domain functions. We confirm this by showing the data from exact diagonalization (ED), black solid line, for $L=1600$ and $\gamma=0.01$, where the oscillation show a regular decay while as there is sudden disappearance of oscillations from the corresponding plots from the numerical Laplace inverse.
Note that initially the brown dots overlap with the red crosses, as the initial dynamics will not show any finite size effects. However, for $t\sim L/4$ for the red crosses, they start to sharply go below $\beta=0.5$ as the finite size effect kick in, on the other hand the brown dots saturate to $\beta=0.5$. The factor of $1/4$ in the time scale for the finite size effects arises due to periodic boundaries as the two fronts in the domain wall at $x=L/2$ and $x=L$ move towards each other and meet when $t\sim L/4$.
Let us now consider the off-diagonal elements of the correlation matrix. Their solution in the limit $L\rightarrow\infty$ follows from using Eq. (37) into Eq. (35). Doing so and replacing $\sin(zl)$ by $(e^{izl}-e^{-izl})/(2i)$ we get,
$\mathcal{G}_{l,0}=\frac{e^{izl}}{2\omega\sin z}\bigg{(}1+\frac{i\omega e^{-iz}-s}{\sqrt{\tilde{s}^{2}+\omega^{2}}-4\gamma}\bigg{)},$ (46)
@@
The asymptotic behaviors for long and short times is obtained by looking at the small $s\sim\omega(q)>4\gamma$ and large $s\sim\omega(q)\sim q<4\gamma$ behavior of $\mathcal{G}_{l,0}$, respectively. We once again choose the initial condition to be $c_{x}=\delta_{x,0}$. The small $s$ behavior is given by,
@@
$\mathcal{G}_{l,0}^{d}\approx\left(\frac{i}{2\gamma}\right)^{l}\frac{q^{l}}{s+2q^{2}/\gamma}$ (50)
@@
$C_{x+l,x}(t)\approx\left(\frac{i}{2\gamma}\right)^{l}\frac{\partial^{l}}{\partial x^{l}}\frac{e^{-\frac{(x-l/2)^{2}}{8t/\gamma}}}{\sqrt{8\pi t/\gamma}}.$ (51)
Therefore, we can see that the for $|x-l/2|<8t/\gamma$, the leading behavior of $|C_{x+l,x}|\sim\frac{1}{(\gamma t)^{k+1/2}}$ for $l=2k-1$ and $l=2k$, where $k=1,2,3,...$. In Fig. 4, we show that agreement of these behaviors with the exact numerical computation.

### v2

@@ III.1 Other Models
$\begin{pmatrix}a_{1}(q)&&a_{2}(q)&&a_{3}(q)&&a_{4}(q)\\ 1&&0&&0&&0\\ 0&&1&&0&&0\\ 0&&0&&1&&0\\ \end{pmatrix},$ (34)
@@
Once the transfer matrix is known, behavior of the product can be studied numerically for large and small $s$ limit for thermodynamically large system sizes, which as we will see in the next section determines the asymptotic behavior of the correlators at short and long times, respectively.
III.2 Off-diagonal elements and non-diagonal initial states
We now obtain the solution for the off-diagonal elements of $C(t)$ in terms of $\mathcal{G}_{0,0}$. The $l^{th}$ off-diagonal elements can be written as,
@@ III.2 Off-diagonal elements and non-diagonal initial states
$\mathcal{G}_{l,0}=\frac{-1}{i\omega\sin\alpha}\left[\sin(\alpha l)+(i\omega\sin[\alpha(l-1)]-s\sin\alpha l)\mathcal{G}_{0,0}\right].$ (38)
Note the fact that only the elements of the first column of the matrix $\mathcal{G}$ determine the entire correlation matrix is a direct consequence of choosing the initial correlation matrix to be diagonal. For initial states with non-diagonal correlation matrix, it follows from Eq. (23) that one needs to determine the elements of other columns of the matrix $\mathcal{G}$. For instance, in presence of nonzero currents, i.e. nonzero first off-diagonal of the initial correlation matrix, elements of the second column of $\mathcal{G}$ are also required. These can be determined using the same transfer matrix procedure but now considering the second column of equations from the identity in Eq. (29).
@@ IV Thermodynamic limit and Asymptotic behavior
IV.1 Thermodynamic Limit
In the last section, we showed that the solution for all the correlators is basically determined by a single function $\mathcal{G}_{0,0}$ whose dependence on $L$ is explicitly given in Eq. (32). So, taking the thermodynamic limit is taking $L\rightarrow\infty$ for this function. In this limit, we have $\sin(\alpha L)\approx-\frac{e^{-i\alpha L}}{2i}$ since $\alpha=\arccos(-iu)$ has a positive imaginary part as $u>0$. Substituting $\sin(\alpha L)\approx-\frac{e^{-i\alpha L}}{2i}$ in Eq. (32) $\mathcal{G}_{0,0}$ simplifies to,
@@ IV.1 Thermodynamic Limit
$=\frac{1}{\tilde{s}-4\gamma-i\omega(q)e^{i\alpha}}=\frac{1}{\sqrt{\tilde{s}^{2}+\omega(q)^{2}}-4\gamma},$ (40)
@@
$\alpha=\arccos(-iu)=\frac{\pi}{2}+i\log\big{[}u+\sqrt{u^{2}+1}\big{]}~{}.$ (41)
The solution for diagonal elements in the limit $L\rightarrow\infty$ now reads,
where we have converted the sum over $q_{n}$ into an integral. Note that the thermodynamic limit naturally came out of the solution in Eq. (32), this is one advantage of using this method as opposed to taking the thermodynamic limit via the spectral decomposition of $\mathcal{A}$, i.e. Eq. (16).
For initial states in Eq. (18), ignoring the long range correlations $l\geq 2$ and denoting $C_{x,x}=-m(x)$ and $C_{x+1,x}=ij(x)/(4J)$, the resulting equation for $m(x)$ and $j(x)$ in the continuum limit are the same as Eq. (1) and Eq. (2), respectively, with $v^{2}=8J^{2}$. These lead to the telegrapher’s equation for the magnetization density, $m(x)$, and thus we expect a diffusion constant of $2J^{2}/\gamma$ in the long time limit. This must also follow from expanding $\mathcal{F}(\tilde{s},q)$ from Eq. (40) in the long wavelength limit i.e. around $q=0$. After doing so, we get
@@
which is indeed the Green’s function for the telegrapher’s equation, Eq. (3), with $v^{2}=8J^{2}$, as expected. While the telegrapher’s equation gives the correct diffusive behavior in the long time limit, it does not not give the correct relaxation to the diffusive behavior as it ignores the dispersion, $\omega(q)$, of the system.
IV.2 Asymptotic Behaviors and Analytic structure
Eq. (42) provides a simple and useful expression for the density for following reasons. The long time and short time behavior follow very elegantly by simply looking at this equation in small $s$ and large $s$ limit, respectively. Let us look at these two limits separately and for simplicity we consider $c(q)=1$ i.e. an initial state which gives the magnetization density to be $c_{x}=\delta_{x,0}$.
Short time limit: In this limit we consider $s\sim\omega(q)>4\gamma$, so dropping $4\gamma$ in the denominator of Eq. (42) we get,
@@ IV.2 Asymptotic Behaviors and Analytic structure
where the Bessel function, $J_{0}[\omega(q)t]$ arises from the inverse Laplace of $1/\sqrt{s^{2}+\omega^{2}(q)}$, and the square of the Bessel $J_{x}(4Jt)$, known to be the Green’s function in absence of dephasing Economou (2006), comes by the subsequent integration over $q$. We see that the ballistic behavior starts disappearing exponentially as soon as the dephasing is turned on.
@@
For the long time limit, we consider $4\gamma>s\sim\omega(q)$. Also, since the behavior is dominated by small $q$, i.e. long wavelengths, $\omega(q)\sim 4Jq$. Using this and extending the limits in the $q$ integral from $\infty$- ∞ to $\infty$ as the dominant contribution comes near $q=0$, we have
@@
where $D=2J^{2}/\gamma$ is the diffusion constant which is in accordance with the expectations from the telegrapher’s equation. Note that the small $s$ and $q$ behavior of $\mathcal{G}_{0,0}$ is given by $(s+(2J^{2}q^{2}/\gamma))^{-1}$, and its Laplace inverse gives the diffusion kernel, i.e. $e^{-2J^{2}q^{2}t/\gamma}$, as expected. A deviation from this behavior in $\mathcal{F}(s,q)$ would give rise to anomalous behavior.
@@
The integrand has branch points at $b_{\pm}=\pm i\omega(q)$, and we take the branch cut to run along the imaginary axis between the two branch points. There are also two poles located at $s_{\pm}=\pm\sqrt{16\gamma^{2}-\omega^{2}(q)}$. The poles lie on the real axis for $4\gamma>\omega(q)$ and on the branch cut along imaginary axis for $4\gamma<\omega(q)$. The branch cut joins the two Riemann sheets and the contours of integration(see Fig. 1) for the Laplace inverse are taken on the principal Riemann sheet.
The contribution due to the branch cut is ballistic, and is sensitive to the details of the spectrum of the Hamiltonian i.e. the dispersion given by $\omega(q)$. When the poles lie on the imaginary axis, their contribution is ballistic akin to forward and backward propagating plane waves of frequency $\sqrt{\omega^{2}(q)-16\gamma^{2}}$. However, as $q$ changes such that $\omega(q)$ approaches $4\gamma$ the poles sweep along the imaginary axis towards the origin and meet at the origin at $4\gamma=\omega(q)$. For $4\gamma>\omega(q)$ the poles are off the branch cut and shift to the real axis but on different Riemann sheets joined by the branch cut. Therefore, one has to be careful about which pole contributes. $s_{+}$ shifts to the principal Riemann sheet, and $s_{-}$ shifts to the second Riemann sheet. Thus, only $s_{+}$ contributes and gives a diffusive contribution. This is also evident from Eq. (47) where the contribution from the pole at $s=-2q^{2}/\gamma$, which is just $s_{+}$ at small $q$ shifted by $4\gamma$, gives rise to the diffusion kernel. Physically, the contribution of the pole at $s_{-}$ and $s_{+}$ correspond to evolution backward and forward in time, respectively. Because the diffusion equation (holding in small $q$ or long-wavelength limit) has a meaning only evolving forward in time (going backwards in time any inhomogeneity would get sharper, instead of smoother, violating relaxation) only $s_{+}$ has to be kept. For the ballistic short-wavelength part at $4\gamma<\omega(q)$, where the dephasing scattering length $\sim 1/\gamma$ is larger than the wavelength $\sim 1/q$ and one deals with the wave equation. Therefore, both forward and backward in time evolution are physical and both contributions have to be included.
@@
IV.3 Numerical Results
Eq. (42) can be directly evaluated numerically by summing over the allowed values of $q$ in place of the integral over $q$ and taking the Laplace inverse numerically using standard approaches such as Talbot’s method Dingfelder and Weideman (2015); Cohen (2007). A nice implementation of these methods can be found in mpmath mpmath development team (2023) python library. While for the XX model with dephasing we can use Eq. (42) directly to get the density but in general one needs to multiply the transfer matrices and invert the product to get $\mathcal{G}_{0,0}$ which is then Laplace inverted for every allowed value of $q$. Therefore, assuming that the transfer matrices are of dimension $r$, the complexity of such a calculation scales as $L^{2}r^{6}$. Given that $r$ is of the order of range of couplings which is usually small, the complexity of this method is better than obtaining the density at any time $t$ via exact diagonalization or iterations of matrix Eq. (10) of size $\sim L^{2}$, which have a complexity of about $L^{6}$. Using the transfer matrix approach magnetization profiles for system sizes of the order of $10^{6}$ can be easily obtained which is much bigger than $L\approx 10^{4}$ feasible through diagonalization Varma et al. (2017).
Fig. 2 shows the evolution of the magnetization density, $\langle\sigma_{3}^{x}\rangle=-C_{x,x}(t)$ for $L=10^{5}$ starting from a domain wall initial state using Eq. (42). By domain wall initial state, we mean a product state of Eq. (18) where spins with $x<L/2$ are in the up state and spins with $x>L/2$ are in the down state. The slow down of the propagation of the front is clearly visible from Fig. 2 as $\gamma t$ crosses unit value.
@@ IV.3 Numerical Results
The plot for $\beta(t)$ is shown in panel (b) of Fig. 3. We find that $\beta(t)$ shows oscillations which decay with time, see the brown dots $(L=10^{5},\gamma=0.01)$ and red crosses ($L=1600$, $\gamma=0.01$) in Fig. 3b. The oscillations arise as at small times the Green’s function is given by Bessel functions, and their amplitude goes down as a power-law (nearly ${}1/t$). The bare time scale for the oscillations is given by $t\sim 1/J$, which in the scaled time, $\gamma t$, goes as $\gamma/J$. Therefore, as $\gamma$ is lowered the oscillations shift towards smaller and smaller values of $\gamma t$. We can see this in the blue dashed curve ($L=10^{5}$, $\gamma=0.001$) where the oscillations have already become too small to be captured by Talbot’s numerical Laplace inverse, and one gets a monotonically decreasing curve from ballistic behavior, $\beta=1$, to diffusive behavior, $\beta=0.5$. Talbot’s method fails to capture the oscillations as the oscillation strength becomes small because of its limitations with oscillatory time domain functions. We confirm this by showing the data from exact diagonalization (ED), black solid line, for $L=1600$ and $\gamma=0.01$, where the oscillation show a regular decay while as there is sudden disappearance of oscillations near $\gamma t\approx 10^{-1}$ from the corresponding plot (red crosses) obtained using the numerical Laplace inverse.
Note that initially the brown dots overlap with the red crosses, as the initial dynamics will not show any finite size effects. However, for $t\sim L/4$ for the red crosses, they start to sharply go below $\beta=0.5$ as the finite size effect kicks in, on the other hand the brown dots saturate to $\beta=0.5$. The factor of $1/4$ in the time scale for the finite size effects arises due to periodic boundaries as the two fronts in the domain wall at $x=L/2$ and $x=L$ move towards each other and meet when $t\sim L/4$.
IV.4 Thermodynamic limit of the Off-diagonal Elements
Let us now consider the off-diagonal elements of the correlation matrix. Their solution in the limit $L\rightarrow\infty$ follows from using Eq. (40) in Eq. (38). Doing so and replacing $\sin(\alpha l)$ by $(e^{i\alpha l}-e^{-i\alpha l})/(2i)$ we get,
$\mathcal{G}_{l,0}=\frac{e^{i\alpha l}}{2\omega\sin\alpha}\bigg{(}1+\frac{i\omega e^{-i\alpha}-s}{\sqrt{\tilde{s}^{2}+\omega^{2}}-4\gamma}\bigg{)},$ (49)
@@ IV.4 Thermodynamic limit of the Off-diagonal Elements
The asymptotic behaviors for long and short times is obtained by looking at the small $s\sim\omega(q)<4\gamma$ and at the large $s\sim\omega(q)>4\gamma$ behavior of $\mathcal{G}_{l,0}$, respectively. We once again choose the initial condition to be $c_{x}=\delta_{x,0}$. The small $s$ behavior is given by,
@@
$\mathcal{G}_{l,0}^{d}\approx\left(\frac{i}{2\gamma}\right)^{l}\frac{q^{l}}{s+2J^{2}q^{2}/\gamma}$ (53)
@@
$C_{x+l,x}(t)\approx\left(\frac{i}{2\gamma}\right)^{l}\frac{\partial^{l}}{\partial x^{l}}\frac{e^{-\frac{(x-l/2)^{2}}{8J^{2}t/\gamma}}}{\sqrt{8\pi J^{2}t/\gamma}}.$ (54)
Therefore, for $|x-l/2|<8J^{2}t/\gamma$, the leading behavior of $|C_{x+l,x}|\sim\frac{1}{(\gamma t)^{k+1/2}}$ for $l=2k-1$ and $l=2k$, where $k=1,2,3,...$. In Fig. 4, we show that agreement of these behaviors with the exact numerical computation.